OpenStudy (anonymous):

Mark all ordered pairs which ARE a solution for y < 3 - 2x. A. (0, 4) B. (-1, 3) C. (6, -8) D. (-4, 5)

OpenStudy (anonymous):

B & D

OpenStudy (anonymous):

3 < 3 - 2(-1) = 3 < 3+2 = 3 <5 (correct) 5 < 3 - 2(-4) = 5 < 3+8 = 5 <11 (correct)